For the following recursive functions, Prove that am,n = m · n. Here am,n is defined recursively for (m, n) ∈ N × N.

am.n = { 0 if m=0

n+am-1,n if m!=0


1
Expert's answer
2021-10-28T06:58:24-0400

For "m=0" formula is obviously fulfilled. Check it for "m>0" :

"a_{m,n}=n+a_{m-1,n}=(n+n)+a_{m-2,n}=\ldots=n\cdot m+ a_{0,n}=n\cdot m"